what numbers is 2235, 3864, 450, 72424, 6168, 2421, 432, and 8304 divisible by???? answer back ASAP!! PLEASE
THANKS SOOO MUCHH,
SINCERILY
ariella melamed :)

Answer by checkley77(12844)

factors of 450=2*3*3*5*5
factors of 432=2*2*2*2*3*3*3
therefore it must be 2 or 3
72424/3=2414.333 thus 3 is not the answer.
2235/2=1117.5 thus 2 is not the answer.
thus there is no common fact within this group of numbers except for (1).